25 ml of a solution containg NaOh and Na2CO3 required (I) 50 ml of 0.2 N HCl using phenolphthalein (2) 72 ml of 0.2 N HCl when tirated using methyl Orange as indicator. The ratio of weight of NaOH and Na2CO3 in the solution is
(1) 0.48. (2) 0.44
(3) 0.40. (4) 0.38
1
Expert's answer
2018-04-07T08:12:10-0400
½ Meq. of Na2CO3 + Meq. of NaOH = Meq. of HCl
a/2 + b = 50*0.2 = 10
When MeOH
½ Meq. of Na2CO3 + Meq. of NaOH = Meq. of HCl
a + b = 72*0.2 = 14
Na2CO3 = 14/24
NaOH = 10/24.